# Generative Psycho-Lexical Approach for Constructing Value Systems in Large Language Models

### v1

@@ B.1 Comprehensiveness of the Collected Lexicon
We draw real-world Human-LLM conversations from LMSYS-Chat-1M [95]. Its 1M conversations are clustered into 20 types in the original literature. Some of the types are purely task-oriented, such as discussing software errors and solutions, while others are more likely to involve value-laden interactions, such as role-playing and discussing various characters. We select four types of conversations of the latter category: 1) discussing and describing various characters, 2) creating and improving business strategies and products, 3) discussing toxic behavior across different identities, and 4) generating brief sentences for various job roles. Since the open-source dataset does not provide the cluster labels, we label the conversations using text-embedding-3-large as the zero-shot classifier [58]. After that, we randomly select 512 conversations covering all four types as the test dataset.
We follow GPV [92] to parse the conversations into perceptions and detect the relevant values among 123 values in our lexicon. The results show that each interaction is associated with 47.6 unique values in our lexicon on average, and all interactions are associated with at least 6 unique values. This indicates that the collected lexicon comprehensively captures the values underlying real-world human-LLM interactions. As a contrasting example, Schwartz’s 10 values only cover 93% of all conversations.
@@ B.3 Consistency Across Datasets
In § 6, we use the default profiling prompt for evaluating cross-dataset consistency. Table 11 displays the intra-LLM correlation of the two value measurements for each LLM. In addition, we include the model safety scores from SALAD-Bench [47] for reference. The results show that the intra-LLM correlation is generally high (0.87 on average), indicating the robustness of our value system. We also find that the intra-LLM value consistency is positively correlated with its safety (0.73).
Model Correlation Safety score
gemma-2b 0.953 95.90
@@ B.4 Correlation with Schwartz’s Values
Social Responsibility Risk-Taking Rule-Following Self-Competence Rationality
Self-Direction 0.049 0.595 -0.356 0.164 0.083
@@
We also measure the Schwartz’s values [75] of the LLMs subjects. Table 12 presents the correlation between our factors and Schwartz’s values. The results can offer additional insights into our values.
@@ B.5 Contributions of Values to Safety
Appendix C Comparative Analysis Against ValueLex [9]
Biedma et al. [9] proposed a lexical approach to constructing value systems for LLMs. While their work offers valuable contributions, we believe that core aspects of their approach could benefit from further theoretical and empirical development. In the following, we present a detailed comparative analysis of GPLA in relation to their work.
@@ C.1 Lexicon Collection
Biedma et al. [9] employ direct prompts such as "List the words that most accurately represent your value system" to extract value lexicons. However, direct questioning may not fully capture an LLM’s complete spectrum of values. This work, in contrast, uses indirect, contextually guided questions to elicit a more comprehensive expression of values from LLMs.
In human psychology, self-reported values can be incomplete or skewed due to 1) social desirability bias [63, 45], the tendency for people to self-report values that they believe are more socially acceptable; and 2) unconscious or implicit values [29, 31], where individuals may not recognize some deeply rooted values until certain situations bring them into play. LLM literature also reveals the unreliability of self-report results [21, 66, 92].
Empirically, we examine all the collected lexicons in [9] and find that, exemplified using Schwartz’s values, three values are missing: Achievement, Self-Direction, and Hedonism, according to the embedding similarity criteria established in [82] (cosine similarity < 0.53). Using our indirect, contextually guided questions, we can capture these values, as shown in the following elicited LLM perceptions.
@@ C.2 Computing Value Correlations
The method proposed by Biedma et al. [9] derives correlations based on the co-occurrence of value lexicons in LLM self-reports in response to direct prompts. However, this approach lacks a theoretical foundation in psychology. We hypothesize that the co-occurrence of value lexicons in LLM responses does not necessarily indicate a true correlation between values. To test this hypothesis, we examine the correlation derived from their method for Schwartz’s values.
@@
We pair each of the collected value lexicons in [9] with the most semantically similar Schwartz’s value according to the embedding model [58]. Then, we can compute the correlation between Schwartz’s values using the normalized co-occurrence frequency of the original lexicons, following the method in [9].
@@ Results.
Fig. 5 illustrates the correlation heatmap of Schwartz’s values based on the co-occurrence frequency of the lexicons. The values are ordered along the x and y axes according to Schwartz’s circular structure. Except for the two values of Self-Enhancement, correlations between values generally contradict the theoretical structure of Schwartz’s values. The results suggest that the co-occurrence frequency of lexicons in responses to direct prompts does not necessarily reflect the true value correlation. Value measurements using GPV are more theoretically grounded and empirically validated [92].
Figure 5: Correlation heatmap derived from lexicon co-occurrence frequency [9], after Min-Max normalization.
@@ C.3 LLM Subjects
Biedma et al. [9] tune the generation hyperparameters (e.g., temperature, top-p) for different LLMs, attempting to generate diverse LLM subjects. However, simply tuning the generation hyperparameters is not sufficient to ensure the diversity and coverage of the LLM subjects, as they are not effective in steering the LLMs toward certain value orientations [67]. In this work, we employ the validated value-anchoring prompts [67] to steer the LLMs toward specific value orientations, ensuring the best possible coverage of the value spectrum and the practical relevance of our measurement results, since steering LLMs toward certain roles is common in public-facing applications nowadays.

### v2

@@ B.1 Comprehensiveness of the Collected Lexicon
We draw real-world Human-LLM conversations from LMSYS-Chat-1M zheng2023lmsyschat1m. Its 1M conversations are clustered into 20 types in the original literature. Some of the types are purely task-oriented, such as discussing software errors and solutions, while others are more likely to involve value-laden interactions, such as role-playing and discussing various characters. We select four types of conversations of the latter category: 1) discussing and describing various characters, 2) creating and improving business strategies and products, 3) discussing toxic behavior across different identities, and 4) generating brief sentences for various job roles. Since the open-source dataset does not provide the cluster labels, we label the conversations using text-embedding-3-large as the zero-shot classifier openai2024textembedding3large. After that, we randomly select 512 conversations covering all four types as the test dataset.
We follow GPV ye2025gpv to parse the conversations into perceptions and detect the relevant values among 123 values in our lexicon. The results show that each interaction is associated with 47.6 unique values in our lexicon on average, and all interactions are associated with at least 6 unique values. This indicates that the collected lexicon comprehensively captures the values underlying real-world human-LLM interactions. As a contrasting example, Schwartz’s 10 values only cover 93% of all conversations.
@@ B.3 Consistency Across Datasets
In § 6, we use the default profiling prompt for evaluating cross-dataset consistency. Table 10 displays the intra-LLM correlation of the two value measurements for each LLM. In addition, we include the model safety scores from SALAD-Bench li2024salad for reference. The results show that the intra-LLM correlation is generally high (0.87 on average), indicating the robustness of our value system. We also find that the intra-LLM value consistency is positively correlated with its safety (0.73).
[…]
@@ B.4 Correlation with Schwartz’s Values
[…]
@@
We also measure the Schwartz’s values schwartz1992universals of the LLMs subjects. Table 11 presents the correlation between our factors and Schwartz’s values. The results can offer additional insights into our values.
@@ B.5 Contributions of Values to Safety
Appendix C Comparative Analysis Against ValueLex (biedma2024beyond)
biedma2024beyond proposed a lexical approach to constructing value systems for LLMs. While their work offers valuable contributions, we believe that core aspects of their approach could benefit from further theoretical and empirical development. In the following, we present a detailed comparative analysis of GPLA in relation to their work.
@@ C.1 Lexicon Collection
biedma2024beyond employ direct prompts such as "List the words that most accurately represent your value system" to extract value lexicons. However, direct questioning may not fully capture an LLM’s complete spectrum of values. This work, in contrast, uses indirect, contextually guided questions to elicit a more comprehensive expression of values from LLMs.
In human psychology, self-reported values can be incomplete or skewed due to 1) social desirability bias randall1993social; larson2019controlling, the tendency for people to self-report values that they believe are more socially acceptable; and 2) unconscious or implicit values greenwald1995implicit; hofer2006congruence, where individuals may not recognize some deeply rooted values until certain situations bring them into play. LLM literature also reveals the unreliability of self-report results dominguez2023questioning; rottger2024political; ye2025gpv.
Empirically, we examine all the collected lexicons in biedma2024beyond and find that, exemplified using Schwartz’s values, three values are missing: Achievement, Self-Direction, and Hedonism, according to the embedding similarity criteria established in sorensen2024value (cosine similarity < 0.53). Using our indirect, contextually guided questions, we can capture these values, as shown in the following elicited LLM perceptions.
  • [leftmargin=2em]
@@ C.2 Computing Value Correlations
The method proposed by biedma2024beyond derives correlations based on the co-occurrence of value lexicons in LLM self-reports in response to direct prompts. However, this approach lacks a theoretical foundation in psychology. We hypothesize that the co-occurrence of value lexicons in LLM responses does not necessarily indicate a true correlation between values. To test this hypothesis, we examine the correlation derived from their method for Schwartz’s values.
@@
We pair each of the collected value lexicons in biedma2024beyond with the most semantically similar Schwartz’s value according to the embedding model openai2024textembedding3large. Then, we can compute the correlation between Schwartz’s values using the normalized co-occurrence frequency of the original lexicons, following the method in biedma2024beyond.
@@ Results.
Fig. 5 illustrates the correlation heatmap of Schwartz’s values based on the co-occurrence frequency of the lexicons. The values are ordered along the x and y axes according to Schwartz’s circular structure. Except for the two values of Self-Enhancement, correlations between values generally contradict the theoretical structure of Schwartz’s values. The results suggest that the co-occurrence frequency of lexicons in responses to direct prompts does not necessarily reflect the true value correlation. Value measurements using GPV are more theoretically grounded and empirically validated ye2025gpv.
Figure 5: Correlation heatmap derived from lexicon co-occurrence frequency biedma2024beyond, after Min-Max normalization.
@@ C.3 LLM Subjects
biedma2024beyond tune the generation hyperparameters (e.g., temperature, top-p) for different LLMs, attempting to generate diverse LLM subjects. However, simply tuning the generation hyperparameters is not sufficient to ensure the diversity and coverage of the LLM subjects, as they are not effective in steering the LLMs toward certain value orientations rozen2024llms. In this work, we employ the validated value-anchoring prompts rozen2024llms to steer the LLMs toward specific value orientations, ensuring the best possible coverage of the value spectrum and the practical relevance of our measurement results, since steering LLMs toward certain roles is common in public-facing applications nowadays.